# 2cm2.5cm2.5cm2cm

### v2

@@ Abstract
We propose a general class of INteger-valued Generalized AutoRegressive Conditionally Heteroscedastic (INGARCH) processes by allowing time-varying mean and dispersion parameters, which we call time-varying dispersion INGARCH (tv-DINGARCH) models. More specifically, we consider mixed Poisson INGARCH models and allow for dynamic modeling of the dispersion parameter (as well as the mean), similar to the spirit of the ordinary GARCH models. We derive conditions to obtain first and second-order stationarity, and ergodicity as well. Estimation of the parameters is addressed and their associated asymptotic properties are established as well. A restricted bootstrap procedure is proposed for testing constant dispersion against time-varying dispersion. Monte Carlo simulation studies are presented for checking point estimation, standard errors, and the performance of the restricted bootstrap approach. We apply the tv-DINGARCH process to model the weekly number of reported measles infections in North Rhine-Westphalia, Germany, from January 2001 to May 2013, and compare its performance to the ordinary INGARCH approach.
@@ 1 Introduction
Modeling count time series data is a challenging and very exciting research topic with applications in many different areas such as epidemiology, sociology, economics, and health science. A well-established methodology, for dealing with count time series data, has been developed under the framework of INteger-valued Generalized AutoRegressive Conditional Heteroscedastic (INGARCH) models, which have been initially studied and explored by Heinen (2003), Ferland et al. (2006), Fokianos et al. (2009), and Fokianos (2011). The INGARCH nomenclature emerges from the fact that for a Poisson distribution (this assumption is imposed in the aforementioned papers), the mean equals its variance, therefore modeling of mean implies modeling of variance like the classic continuous GARCH models introduced by Bollerslev (1986). Consequently, such models are considered in the literature as integer counterpart of the GARCH models although this terminology should be used cautiously because ordinary GARCH processes do not consider dynamics for the conditional mean; for instance, see Fokianos et al. (2009).
A Poisson $\mbox{INGARCH}(p,q)$ (with $p,q\in\mathbb{N}_{0}$) model $\{Y_{t}\}_{t\in\mathbb{N}}$ is defined by
@@
where $\mathcal{F}_{t-1}=\sigma\{Y_{t-1},\ldots,Y_{0},\lambda_{0}\}$, $\beta_{0}>0$, $\beta_{i}\geq 0$ and $\alpha_{j}\geq 0$, for all $i=1,\ldots,p$ and $j=1,\ldots,q$. Since $E(Y_{t}|\mathcal{F}_{t-1})=\mbox{Var}(Y_{t}|\mathcal{F}_{t-1})=\lambda_{t}$, the model specifies dynamic processes for both the conditional mean and variance, as previously stated.
@@
The Poisson INGARCH model has been extensively considered in the literature. However, due to limitations for fitting adequately real count time series data (e.g. not capturing all the sources of overdispersion although the model itself is overdispersed, inflation or deflation of zeros), other variants have been proposed based on such as the negative binomial INGARCH process by Zhu (2011) and Christou and Fokianos (2014) or the more general mixed Poisson INGARCH models by Christou and Fokianos (2015), Silva and Barreto-Souza (2019), among others. Zero-inflated versions of the Poisson and negative binomial INGARCH models were proposed by Zhu (2012a), while processes dealing with both overdispersion and underdispersion were proposed by Zhu (2012b, c) and Xu et al. (2012).
@@
An alternative process to the linear model for the conditional mean, as in Equation (1), was introduced by Fokianos and Tjøstheim (2011) through the log-linear INGARCH processes. These models cope with both negative and positive autocorrelation function and allow inclusion of covariates in a straightforward way. Another different approach was recently proposed by Weiß et al. (2022), where a softplus link function is assumed instead of a logarithmic one.
@@
Based on the discussion so far, the Poisson and Negative Binomial distributions are frequently used in applications. In several cases, the Negative Binomial model fits more adequate due to the fact that this distribution depends on a dispersion parameter which facilitates flexible modeling. The dispersion parameter is usually assumed fixed and is estimated by using the Pearson residuals; see Christou and Fokianos (2014) for more. The main goal of this paper is to propose a novel and general class of INGARCH processes based on the mixed Poisson distributions (the Negative Binomial distribution falls in this class) by allowing time-varying mean and dispersion parameters. We call these models time-varying dispersion INGARCH (tv-DINGARCH) models. Hence, we develop methodology and study new models for integer-valued time series, where the assumption of constant dispersion might not be hold. The advantage of the tv-DINGARCH processes over the ordinary INGARCH models will be illustrated in Section 5, where we show that such processes fit better real data whose stylistic facts cannot be dealt with the existing INGARCH models.
@@
For example, in a count regression context, Barreto-Souza and Simas (2016) demonstrated through a data application on the attendance behavior of high school juniors that a constant dispersion assumption can be violated. Generalized linear models allowing regression structures for both mean and dispersion/precision parameters have been discussed previously by Efron (1986) and Smyth (1989), among others. In this work, we develop further this line of research by studying in detail joint modeling of mean and dispersion parameters in the context of dependent data. We put special emphasis in the case of INGARCH models for count time series and demonstrate the usefulness of such methodological development. Besides exploring the traditional INGARCH models, the proposed class contains the case that admits constant mean and time-dependent variance, similar to the spirit of the ordinary GARCH models. Another important feature is volatility modeling, which is a well-explored topic for the case of continuous valued time series but has been neglected in the context of integer-valued time series. Although the traditional Poisson INGARCH processes consider a time-dependent conditional variance, this in turn is driven by the dynamics of the mean process. The main disadvantage of this approach is that imposes severe restrictions as illustrated in the real data application in this paper. The tv-DINGARCH models relax such an assumption by considering a time-dependent dispersion parameter, therefore also controlling the conditional variance and allowing for additional source of volatility.
@@
In the context of time-varying models for INGARCH processes, Roy and Karmakar (2021) introduced a Poisson INGARCH model with time-varying coefficients and Ratnayake and Samaranayake (2023) study a time-varying INGARCH model based on the zero-inflated Poisson distribution. These authors consider dynamics for the inflation parameter which is assumed to depend on exogenous variables. Furthermore, Doukhan et al. (2022) study in detail first order non-stationary INGARCH models and prove mixing conditions under natural assumptions. Our approach is different in several aspects. First, we consider mixed Poisson distributions (rather than Poisson and zero-inflated Poisson models) which cover a broad class of count distributions that include the Poisson and Negative Binomial. In addition, we consider dynamics that enter through the dispersion parameter. One advantage of the latter feature is that we are able to establish desirable stability properties for the models we consider, such as stationarity and ergodicity, based on e-chain theory.
@@
The paper is organized as follows. In Section 2, we define the class of tv-DINGARCH models and then derive its stochastic properties. We establish conditions ensuring stationarity and ergodicity of the count processes. Section 3 is devoted to the estimation of the parameters and the derivation of their associated asymptotic properties as well. Aiming at testing constant dispersion in practice, a restricted bootstrap procedure is proposed in Section 4. Monte Carlo simulation studies are presented for checking point estimation, standard errors, and the performance of the restricted bootstrap approach. In Section 5, we apply the tv-DINGARCH process to model the weekly number of reported measles infections in North Rhine-Westphalia, Germany, from January 2001 to May 2013, and compare its performance to the ordinary INGARCH approach. Some technical results and proofs are provided in the Appendix. This paper contains Supplementary Material, which is available upon request.
@@
We say that a random variable $Y$ follows a mixed Poisson (MP) distribution if satisfies the stochastic representation $Y|Z=z\sim\mbox{Poisson}(\lambda z)$, with $Z$ following some non-negative distribution with $E(Z)=1$ (standardization) and $\mbox{Var}(Z)=\phi^{-1}$, for $\lambda,\phi>0$. We denote $Y\sim\mbox{MP}(\lambda,\phi)$. In this case, $E(Y)=\lambda$ and $\mbox{Var}(Y)=\lambda+\lambda^{2}/\phi$. A random variable $Z$ following a Gamma distribution, with respective shape and scale parameters $a>0$ and $b>0$, is denoted by $Z\sim\mbox{Gamma}(a,b)$, where $E(Z)=a/b$ and $\mbox{Var}(Z)=a/b^{2}$. If the mixing variable $Z\sim\mbox{Gamma}(\phi,\phi)$, then we obtain that $Y$ follows a negative binomial (NB) distribution, i.e. $Y\sim\mbox{NB}(\lambda,\phi)$.
For an $l$-dimensional vector ${\bf x}=(x_{1},\ldots,x_{l})^{\top}$, let $\|{\bf x}\|_{p}=(\sum_{i=1}^{l}|x_{i}|^{p})^{1/p}$, for $p\in[1,\infty)$, and $\|{\bf x}\|_{p}=\max_{1\leq i\leq l}|x_{i}|$ for $p=\infty$. The induced $p$-norm for a $m\times l$ matrix ${\bf C}$ is then defined by $\|{\bf C}\|_{p}=\max_{{\bf x}\neq{\bf 0}}\{\|{\bf Cx}\|_{p}/\|{\bf x}\|_{p}:\,{\bf x}\in\mathbb{R}^{l}\}$, for $p\in[1,\infty]$.
@@ 2 The tv-DINGARCH Models
In this section we propose a class of time-varying dispersion INGARCH models by allowing both the mean and the dispersion parameter to depend on lagged valued of the observed series and their past values as follows.
@@
where $\mathcal{F}_{t-1}=\sigma\{Y_{t-1},\ldots,Y_{0},\phi_{0},\lambda_{0}\}$, and $(\lambda_{0},\phi_{0})$ denoting some starting value, $f:\mathbb{N}^{p_{1}}\times(0,\infty)^{q_{1}}\rightarrow(0,\infty)$, and $g:\mathbb{N}^{p_{2}}\times(0,\infty)^{q_{2}}\rightarrow(0,\infty)$.
@@
The negative binomial model will be assumed to prove finiteness of moments and to derive the asymptotic theory for conditional maximum likelihood estimators in what follows. Some of results hold in general for mixed Poisson models, see Theorem 2.1, which states stationarity and ergodicity of time-varying dispersion INGARCH models.
Definition 2.1 is general and some additional assumptions on the function $f$ and $g$ are necessary to study in detail these processes. We consider the first order model ($p_{1}=p_{2}=q_{1}=q_{2}=1$) linear tv-DINGARCH processes, which will be defined in what follows. This particular linear parametric form is a common choice considered in the literature but also enables a thorough study of stability properties of processes defined by Definition 2.1.
@@
Note that we are using $Y_{t-1}$ to model $\phi_{t}$ in eq. (3), and more generally in Def. 2.1, instead of terms such as $Y^{2}_{t-1}$ and $|Y_{t-1}|$ considered for the continuous GARCH models and their modification. This is so because $Y_{t-1}$ is non-negative in our case (count time series).
@@
Some particular cases are obtained from the linear tv-DINGARCH class defined above. For $\alpha_{1}=\alpha_{2}=0$, we obtain the ordinary linear mixed Poisson INGARCH (Christou and Fokianos, 2015; Silva and Barreto-Souza, 2019) models as particular cases. Additionally, by taking $\alpha_{0}\rightarrow 0^{+}$, we also obtain the Poisson INGARCH model as a limiting member of our proposed class. Another interesting and novel model arises when $\beta_{1}=\beta_{2}=0$. Under this setting, the mean of the INGARCH process is constant and the variance is time-dependent as in the case of ordinary GARCH models (Bollerslev, 1986). Such property does not hold for the standard Poisson INGARCH model (1). In this case, we refer to (3) as the Pure INGARCH (P-INGARCH) process. We use the term “pure" to connect the fact that our model mimics the traditional continuous GARCH models (constant mean and time-varying variance). Hence, our approach is general and encompasses many different models studied earlier in the literature.
Some simulated trajectories of the linear tv-DINGARCH models for some parameter settings are shown in the Supplementary Material.
@@ 2.1 Stationarity and Ergodicity
Conditions for the existence and stationarity of the process (3) can be established, for example, by following the strategy by Christou and Fokianos (2014), which relies on establishing weak dependence by Doukhan and Wintenberger (2008). We here obtain such desirable properties based on another approach, the e-chains theory (Meyn and Tweedie, 1993), as follows.
The dynamic latent processes $\{\lambda_{t}\}_{t\geq 1}$ and $\{\phi_{t}\}_{t\geq 1}$ given in (3) can be rewritten in a matrix form. By defining $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}$, for $t\geq 1$, $\boldsymbol{\tau}=(\beta_{0},\alpha_{0})^{\top}$ ${\bf A}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix}$, and ${\bf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}$, we have that $\boldsymbol{\xi}_{t}=\boldsymbol{\tau}+{\bf B}(Y_{t-1},Y_{t-1})^{\top}+{\bf A}\boldsymbol{\xi}_{t-1}$. Under this matrix representation, we can adapt the strategy in Liu (2012) to establish the existence and stationarity of our process. In that work, the author provided stochastic properties of a bivariate Poisson INGARCH model. The key point is to show that $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{N}}$ is an e-chain Meyn and Tweedie (1993, Ch.18). The following theorem holds true for any mixed Poisson model including the Negative Binomial model.
@@
Let $\{Y_{t}\}_{t\in\mathbb{N}}$ be a tv-DINGARCH process as in (3). If there exists $p\in[1,\infty]$ such that $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$, then the trivariate process $\{(Y_{t},\lambda_{t},\phi_{t})\}_{t\in\mathbb{N}}$ has a unique stationary and ergodic solution.
@@
Then by considering $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}=\boldsymbol{\tau}+\sum_{j=1}^{q}{\bf B}_{j}(Y_{t-j},Y_{t-j})^{\top}+\sum_{i=1}^{m}{\bf A}_{i}\boldsymbol{\xi}_{t-i}$, where $\{{\bf A}_{i}\}_{i=1}^{m}$ and $\{{\bf B}_{j}\}_{j=1}^{q}$ are
@@
$m=\max(p_{1},p_{2})$, and $q=\max(q_{1},q_{2})$. Then, following the same steps as in the proof of Theorem 2.1 the following result can be established.
@@
We close this section by proving finiteness of the moments (see Appendix for a proof) of the first-order negative binomial INGARCH process, which will be used in the next section to establish the asymptotic normality of the conditional maximum likelihood estimators.
@@
For $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, the first-order negative binomial INGARCH process $\{Y_{t}\}_{t\in\mathbb{N}}$ given by (3) has finite moments of order $k\in\mathbb{N}$, for an arbitrary $k$.

### v3

@@ Abstract
We introduce a general class of INteger-valued Generalized AutoRegressive Conditionally Heteroscedastic (INGARCH) processes by allowing simultaneously time-varying mean and dispersion parameters. We call such models time-varying dispersion INGARCH (tv-DINGARCH) models. More specifically, we consider mixed Poisson INGARCH models and allow for dynamic modeling of both mean and dispersion parameters. We derive conditions to obtain stochastic stability of tv-DINGARCH processes. Additionally, we study maximum likelihood estimation in detail including its asymptotic distribution. A restricted bootstrap procedure is proposed for testing constant dispersion against time-varying dispersion. Monte Carlo simulation studies are presented for checking point estimation, standard errors, and the performance of the restricted bootstrap approach. We apply the tv-DINGARCH process to model the weekly number of reported measles infections in North Rhine-Westphalia, Germany, from January 2001 to May 2013, and compare its performance to the ordinary INGARCH approach.
@@ 1 Introduction
Modeling count time series data is a challenging and exciting research topic with applications in many areas such as epidemiology, sociology, economics, and health science. A well-established methodology, for dealing with count time series data, has been developed under the framework of the so-called INteger-valued Generalized AutoRegressive Conditional Heteroscedastic (INGARCH) models, which have been initially studied and explored by Heinen (2003), Ferland et al. (2006), Fokianos et al. (2009), and Fokianos (2011). The INGARCH nomenclature emerges from the fact that for a Poisson distribution (this assumption is imposed in the aforementioned papers), the mean equals its variance. Therefore modeling of mean implies modeling of variance like in the classic GARCH models introduced by Bollerslev (1986). Consequently, such models are considered in the literature as an integer counterpart of the GARCH models although this terminology should be used cautiously because ordinary GARCH processes are employed for modeling time-varying variances.
A Poisson $\mbox{INGARCH}(p,q)$ (with $p,q\in\mathbb{N}$) model $\{Y_{t}\}_{t\in\mathbb{Z}}$ is defined by
@@
where $\mathcal{F}_{t}=\sigma\{Y_{s},s\leq t\}$, $\beta_{0}>0$, $\beta_{i}\geq 0$ and $\alpha_{j}\geq 0$, for all $i=1,\ldots,p$ and $j=1,\ldots,q$. Since $E(Y_{t}|\mathcal{F}_{t-1})=\mbox{Var}(Y_{t}|\mathcal{F}_{t-1})=\lambda_{t}$, the model specifies dynamic processes for both the conditional mean and variance, as previously stated.
@@
The Poisson INGARCH model has been extensively considered in the literature. However, due to limitations for fitting adequately real count time series data (e.g. not capturing all the sources of overdispersion although the model itself is overdispersed), several variants of the Poisson INGARCH model have been proposed. These include the negative binomial (NB) INGARCH process by Zhu (2011) and Christou and Fokianos (2014) or the more general mixed Poisson INGARCH models by Christou and Fokianos (2015), Silva and Barreto-Souza (2019), among others. Zero-inflated versions of the Poisson and negative binomial INGARCH models have been studied by Zhu (2012a), while processes dealing with both overdispersion and underdispersion were proposed by Zhu (2012b, c) and Xu et al. (2012). A negative binomial quasi-likelihood based inference for general integer-valued time series models with application to the Poisson and NB INGARCH models has been developed by Aknouche et al. (2018).
@@
An alternative process to the linear model given in (1), was introduced by Fokianos and Tjøstheim (2011) through the log-linear INGARCH processes. These models cope with both negative and positive autocorrelation functions and allow the inclusion of covariates in a straightforward way. More recently Weiß et al. (2022) considered a softplus link function instead of a logarithmic one.
@@
Based on the discussion so far, the Poisson and negative binomial distributions are frequently used in applications for count time series analysis. In several cases, the negative binomial model fits more adequately because this distribution depends on a dispersion parameter which facilitates flexible modeling. The dispersion parameter is usually assumed fixed and is estimated by using the Pearson residuals; see Christou and Fokianos (2014) for more. The main goal of this paper is to propose a novel and general class of INGARCH processes based on the mixed Poisson distributions (the negative binomial distribution belongs to this class) by allowing both time-varying mean and dispersion parameters. We call these models time-varying dispersion INGARCH (tv-DINGARCH) models. Hence, we develop methodology and introduce new models for integer-valued time series, where the assumption of constant dispersion might not hold. The usefulness of the tv-DINGARCH processes over the ordinary INGARCH models will be illustrated in Section 5, where we show that such processes fit better real data whose stylistic facts cannot be dealt with the existing INGARCH models.
@@
For example, in a count regression context, Barreto-Souza and Simas (2016) demonstrated through a data application on the attendance behavior of high school juniors that a constant dispersion assumption can be violated. Generalized linear models allowing regression structures for both mean and dispersion/precision parameters have been discussed previously by Efron (1986) and Smyth (1989), among others. In this work, we further develop this line of research by studying in detail the joint modeling of mean and dispersion parameters in the context of dependent data. We put special emphasis on the case of INGARCH models for count time series and demonstrate the usefulness of such methodological development. Besides extending the traditional INGARCH models, the proposed class includes the case of constant mean and time-dependent variance, which is a new model class that does not fall in the ordinary INGARCH model approach. Another important feature is volatility modeling, which is a well-explored topic for the case of continuous-valued time series but has been neglected in the context of integer-valued time series. Although the traditional Poisson INGARCH processes consider a time-dependent conditional variance, this in turn is solely driven by the dynamics of the mean process. The main disadvantage of this approach is that imposes severe restrictions on the volatility modelling as illustrated in the real data application in this paper. The class tv-DINGARCH models relaxes this assumption by considering a time-dependent dispersion parameter, therefore also controlling the conditional variance and allowing for additional sources of volatility.
@@
In the context of time-varying models for INGARCH processes, Roy and Karmakar (2021) introduced a Poisson INGARCH model with time-varying coefficients and Ratnayake and Samaranayake (2023) study a time-varying INGARCH model based on the zero-inflated Poisson distribution. These authors consider dynamics for the inflation parameter which is assumed to depend on exogenous variables. Furthermore, Doukhan et al. (2022) study in detail first order non-stationary INGARCH models and prove mixing conditions under natural assumptions. Our approach is different in several aspects. First, we consider mixed Poisson distributions (rather than Poisson and zero-inflated Poisson models) which cover a broad class of count distributions that include the Poisson and negative binomial. In addition, we consider dynamics that enter through the dispersion parameter. One advantage of the latter feature is that we are able to establish desirable stability properties for the models we consider, such as stationarity and ergodicity, by employing e-chain theory; see Meyn and Tweedie (1993).
@@
The paper is organized as follows. In Section 2, we define the class of tv-DINGARCH models and then derive its stochastic properties. We establish conditions ensuring stationarity and ergodicity of the count processes. Section 3 is devoted to the estimation of the parameters and the derivation of their associated asymptotic properties as well. Aiming at testing constant dispersion in practice, a restricted bootstrap procedure is proposed in Section 4. Monte Carlo simulation studies are presented for checking point estimation, standard errors, and the performance of the restricted bootstrap approach. In Section 5, we apply the tv-DINGARCH process to model the weekly number of reported measles infections in North Rhine-Westphalia, Germany, from January 2001 to May 2013, and compare its performance to the ordinary INGARCH approach. Some technical results and proofs are provided in the Appendix. This paper contains Supplementary Material.
@@
We say that a random variable $Y$ follows a mixed Poisson (MP) distribution if satisfies the stochastic representation $Y|Z=z\sim\mbox{Poisson}(\lambda z)$, with $Z$ following some non-negative distribution with $E(Z)=1$ (standardization) and $\mbox{Var}(Z)=\phi^{-1}$, for $\lambda,\phi>0$. In this case, $E(Y)=\lambda$ and $\mbox{Var}(Y)=\lambda+\lambda^{2}/\phi$. We denote $Y\sim\mbox{MP}(\lambda,\phi;\mathcal{D})$, where $\mathcal{D}$ denotes the distribution of $Z$. A random variable $Z$ following a Gamma distribution, with respective shape and scale parameters $a>0$ and $b>0$, is denoted by $Z\sim\mbox{Gamma}(a,b)$, where $E(Z)=a/b$ and $\mbox{Var}(Z)=a/b^{2}$. If the mixing variable $Z\sim\mbox{Gamma}(\phi,\phi)$ (i.e. $\mathcal{D}=\mbox{Gamma}$), then we obtain that $Y$ follows a negative binomial (NB) distribution, i.e. $Y\sim\mbox{NB}(\lambda,\phi)$.
For an $l$-dimensional vector ${\bf x}=(x_{1},\ldots,x_{l})^{\top}$, let $\|{\bf x}\|_{p}=(\sum_{i=1}^{l}|x_{i}|^{p})^{1/p}$, for $p\in[1,\infty)$, and $\|{\bf x}\|_{p}=\max_{1\leq i\leq l}|x_{i}|$ for $p=\infty$. The induced $p$-norm for a $m\times l$ matrix ${\bf C}$ is then defined by $\|{\bf C}\|_{p}=\max_{{\bf x}\neq{\bf 0}}\{\|{\bf Cx}\|_{p}/\|{\bf x}\|_{p}:\,{\bf x}\in\mathbb{R}^{l}\}$, for $p\in[1,\infty]$. For a square matrix A, we denote its spectral radius by $\rho({\bf A})$.
@@ 2 The tv-DINGARCH Models
In this section, we propose a class of time-varying dispersion INGARCH models by allowing both mean and dispersion parameters to depend on the lagged values of the observed series and their past values as follows.
@@
where $\mathcal{F}_{t}=\sigma\{Y_{s},s\leq t\}$ and $f:\mathbb{N}^{p_{1}}\times(0,\infty)^{q_{1}}\rightarrow(0,\infty)$, and $g:\mathbb{N}^{p_{2}}\times(0,\infty)^{q_{2}}\rightarrow(0,\infty)$.
@@
The negative binomial model will be assumed to prove the finiteness of moments and to derive the asymptotic theory for conditional maximum likelihood estimators in what follows.
Definition 2.1 is general and some additional assumptions on the functions $f$ and $g$ are necessary to study in detail these processes. We consider the first-order model ($p_{1}=p_{2}=q_{1}=q_{2}=1$) linear tv-DINGARCH processes, which will be defined in what follows. This particular linear parametric form is a common choice considered in the literature but also enables a thorough study of stability properties of processes given by Definition 2.1.
@@
We use $Y_{t-1}$ to model $\phi_{t}$ in eq. (3), and more generally in Def. 2.1, instead of terms such as $Y^{2}_{t-1}$ and $|Y_{t-1}|$ considered for the continuous GARCH models and their modification. This is so because $Y_{t-1}$ is non-negative as we deal with count data.
@@
Some particular cases are obtained from the tv-DINGARCH class defined above. When $\alpha_{1}=\alpha_{2}=0$, we obtain the ordinary linear mixed Poisson INGARCH (Christou and Fokianos, 2015; Silva and Barreto-Souza, 2019) models as particular cases. Additionally, by taking $\alpha_{0}\rightarrow 0^{+}$, we also obtain the Poisson INGARCH model as a limiting member of our proposed class. Another interesting and novel model arises when $\beta_{1}=\beta_{2}=0$. Under this setting, the mean of the INGARCH process is constant and the variance is time-dependent as in the case of ordinary GARCH models (Bollerslev, 1986). Such property does not hold for ordinary Poisson INGARCH models given by Equation (1). When $\beta_{1}=\beta_{2}=0$ we call (3) Pure INGARCH (P-INGARCH) process. We use the term “pure" to emphasize the fact that the model mimics the traditional continuous GARCH models (constant mean and time-varying variance). Hence, our approach is general and encompasses many different models studied earlier in the literature.
Simulated trajectories of the linear tv-DINGARCH models for a few parameter settings are shown in the Supplementary Material.
@@ 2.1 Stationarity and Ergodicity
The dynamic latent processes $\{\lambda_{t}\}_{t\in\mathbb{Z}}$ and $\{\phi_{t}\}_{t\in\mathbb{Z}}$ given in (3) can be written compactly in a vector form. By defining $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}$, $\boldsymbol{\tau}=(\beta_{0},\alpha_{0})^{\top}$, ${\bf A}=\begin{pmatrix}\beta_{2}&0\\ 0&\alpha_{2}\end{pmatrix}$ and ${\bf B}=\begin{pmatrix}\beta_{1}&0\\ 0&\alpha_{1}\end{pmatrix}$, we obtain that $\boldsymbol{\xi}_{t}=\boldsymbol{\tau}+{\bf B}(Y_{t-1},Y_{t-1})^{\top}+{\bf A}\boldsymbol{\xi}_{t-1}$. Using this representation, we adapt the strategy in Liu (2012) to establish existence and stationarity of (3). In that work, the author provided stochastic properties of a bivariate Poisson INGARCH model. The key point is to show that $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$ is an e-chain; see Meyn and Tweedie (1993, Ch.18). In addition, the property of Geometric Moment Contraction (GMC), as studied by Diaconis and Freedman (1999) and Wu and Shao (2004), enables us to prove the desired ergodicity and uniqueness of the stationary distribution of $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$.
Remark 2.3.
In what follows, we prove all theoretical results for the negative binomial linear tv-DINGARCH models as given in 2.2 except Thm. 2.1 and Corollary 2.1 which hold true for any linear tv-DINGARCH mixed Poisson model including the negative binomial model.
@@
Let $\{Y_{t}\}_{t\in\mathbb{Z}}$ be a tv-DINGARCH process (with fixed mixing distribution $\mathcal{D}$) as in (3).
(a) If $\rho({\bf A}+{\bf B})<1$, then there exists at least one stationary distribution to $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$, where $\rho(\cdot)$ denotes the spectral radius. In addition, if $\|{\bf A}\|_{p}<1$ for some $1\leq p\leq\infty$, then the distribution is unique.
(b) If $\|{\bf A}\|_{p}+2^{1-1/p}\|{\bf B}\|_{p}<1$ for some $p\in[1,\infty]$, then $\{\boldsymbol{\xi}_{t}\}_{t\in\mathbb{Z}}$ is a GMC Markov chain with a unique stationary and ergodic distribution.
@@
If we denote by $\boldsymbol{\xi}_{t}=(\lambda_{t},\phi_{t})^{\top}=\boldsymbol{\tau}+\sum_{j=1}^{q}{\bf B}_{j}(Y_{t-j},Y_{t-j})^{\top}+\sum_{i=1}^{m}{\bf A}_{i}\boldsymbol{\xi}_{t-i}$, where $\{{\bf A}_{i}\}_{i=1}^{m}$ and $\{{\bf B}_{j}\}_{j=1}^{q}$ are defined by
@@
$m=\max(p_{1},p_{2})$, and $q=\max(q_{1},q_{2})$, then, following the same steps as in the proof of Theorem 2.1, the next result can be established; see Liu (2012, Prop. 4.3.1)
@@
We close this section by proving the finiteness of the moments (see Appendix for a proof) of the first-order negative binomial tv-DINGARCH process.
@@
If $\|{\bf A}\|_{1}+\|{\bf B}\|_{1}<1$, then the first-order negative binomial tv-DINGARCH process $\{Y_{t}\}_{t\in\mathbb{Z}}$ given by (3) satisfies that $E|Y_{t}|^{k}<\infty$, $E|\lambda_{t}|^{k}<\infty$ and $E|\phi_{t}|^{k}<\infty$ for any $k\in\mathbb{N}$.
Remark 2.4.
We conjecture that Theorem 2.2 can be extended to hold for any linear mixed Poisson tv-DINGARCH model by combining our arguments and the results of Karlis and Xekalaki (2005, Subsection 3.2) concerning moments of mixed Poisson distributions. This is out of the scope of this work since the focus is on the negative binomial model.